5. The electric field caused by a certain point charge has a magnitude of 5.0 Γ— 10^3 𝑁 𝐢 at a distance 0.10m from the charge. What is the magnitude of the charge?

6. Two particles having charges π‘ž1 = 1.00 Γ— 10^βˆ’19𝐢 and π‘ž2 = 2.00 Γ— 10^βˆ’9𝐢 are separated by a distance 2.00m. At what point is the total electric field due to the two charges equal to zero?

7. Create an illustration of the electric field lines of π‘ž1 = 1𝐢 between π‘ž2 and π‘ž3 both having a charge od -2C.

8. A uniform electric field 𝐸 = 8000 𝑁 𝐢 passing through a flat surface with an area of 𝐴 = 10π‘š^2 . Determine the flux if: a.) the electric field direction and the line perpendicular to the surface is pointing in the same direction. b.) the electric field direction and the line perpendicular to the surface is perpendicular. c.) the angle is 45οΏ½


1
Expert's answer
2022-03-16T18:29:14-0400

5.

"q=\frac{Er^2}k=5.6~\mu C."

6.

"x=\frac r{1+\sqrt{\frac{q_2}{q_1}}}=0.83~m."

8a.

0.

8b.

"\Phi=EA=80~kVm."

8c.

"\Phi=EA\cos\alpha=57~kVm."